given lim(as x approaches a) for ((x^2+bx+4b)/(x-a))= -6, find the values of a and b.


1
Expert's answer
2022-04-08T13:18:58-0400

In our case, we have an uncertainty of type 0/0.

It follows from the fact that the denominator tends to zero when a tends to zero, and hence the numerator also tends to zero, otherwise the limit would tend to infinity.

"x^2+bx+4b=0" if "x=a" and due to factor theorem we can write:

"x^2+bx+4b=(x-a)(x-c)" , where "c" is some constant. To make limit equal to -6 we should choose "c=a+6" .

"\displaystyle \lim_{\mathclap{x\to a}} {\frac{x^2+bx+4b}{x-a}}=\lim_{\mathclap{x\to a}} {\frac{(x-a)(x-a-6)}{x-a}}=""\displaystyle\lim_{\mathclap{x\to a}} {(x-a-6)}=-6"

"x^2+bx+4b=(x-a)(x-a-6)"

"x^2+bx+4b=x^2-x(2a+6)+a^2+6a"

Equating coefficients, we get:

"b=-(2a+6)"

"4b=a^2+6a"

Substituting first equation into the second we will obtain:

"-4(2a+6)=a^2+6a"

"a^2+14a+24=0"

Using Vieta's formulas we can write

"a=-12", "b=-(2a+6)=-(2\cdot(-12)+6)=18"

"a=-2", "b=-(2\cdot(-2)+6)=-2"

Answer: "a=-12" , "b=18" or "a=-2", "b=-2" .